OpenStudy (anonymous):

I have a graph and its points are (5,3) and (-4,-6). What is the equation of the line? Give your answer in slope-intercept form.

OpenStudy (igreen):

Slope intercept form is y = mx + b.

OpenStudy (igreen):

Where m = slope and b = y-intercept.

OpenStudy (anonymous):

So how do I find the slope and intercept @iGreen

OpenStudy (igreen):

First you find the slope by using the equation \(m = \dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

How do I know which one should be x2 or x1? Does it matter?

OpenStudy (igreen):

\(y_2 = -6\) \(y_1 = 3\) \(x_2 = -4\) \(x_1 = 5\)

OpenStudy (igreen):

No not really. I just do it in the order it gives you. (5,3) and (-4,-6) So I say x1 is 5, x2 is -4, y1 is 3, and y2 is -6.

OpenStudy (anonymous):

Okay c: That way they are both -9 so is the slope -9 or 1 since the answer was -9/-9?

OpenStudy (igreen):

-9/-9 is 1 so you are correct, the slope is 1.

OpenStudy (anonymous):

Thank you c: So how do I get the intercept?

OpenStudy (igreen):

That gives you y = 1x + b or y = x + b.

OpenStudy (anonymous):

yea so far I got y=1x+b

OpenStudy (igreen):

Just plug any of the points in the equation to find the y-intercept. (5, 3) y = x + b 3 = 5 + b Subtract b to both sides: 3 - b = 5 Subtract 3 to both sides: -b = 2 Multiply -1 to both sides of the equation: b = -2 So your y-intercept is -2, making your equation y = x - 2.

OpenStudy (anonymous):

Thank you so much!!!